OpenStudy (anonymous):

\[\lim_{x \rightarrow \pi+}\sec2x\] is it positive or negative?

OpenStudy (empty):

Any ideas? Where are you stuck?

OpenStudy (tkhunny):

Is the cosine positive or negative in the neighborhood of \(\pi\)?

OpenStudy (anonymous):

I don't know I just cant't grasp the concept.

OpenStudy (anonymous):

this is one part of calculus I am just struggling with

OpenStudy (empty):

Explain what you do understand and I can help you understand and fill in the gaps if you want to learn.

OpenStudy (anonymous):

When its on the graph i can get it but my teacher doesn't want us using a graph

OpenStudy (anonymous):

usually I can solve backward when I don't understand it

OpenStudy (tkhunny):

Come on... \(\cos(\pi) = -1\) \(\cos(2\pi) = +1\) What do we know about \(\sec(2\pi)\)?

OpenStudy (anonymous):

i have class but I will tr again later. hopefully It will help when i log back on

OpenStudy (empty):

@Clare12 Well you are trying to run before you can walk if you don't use graphs. You need to spend more time visualizing roughly what trigonometric graphs look like with several common values such as \(\pi\) or \(\frac{\pi}{2}\) and others.